OpenStudy (shadowlegendx):

Draw a visual representation for 27%

OpenStudy (shadowlegendx):

How would you guys do this?

OpenStudy (shadowlegendx):

There is the drawing 100 slots then filling in 27 of them way, but does anyone else have other ideas?

OpenStudy (luigi0210):

Pie chart maybe? Does it have to be precise?

OpenStudy (shadowlegendx):

Just has to be a model of 27% then I have to explain why I chose that model My teacher is making us do these old worksheets for home weekend homework ._.

OpenStudy (anonymous):

Bar graph or pie chart works. Know the percent equation?

OpenStudy (luigi0210):

Well if you're allowed to get creative, use a chart sales thing.. like a bacon sales chart thing ;) ~here comes batman~

OpenStudy (luigi0210):

Cause he loves bacon xD

OpenStudy (anonymous):

Luigi, you know I usually wait for you to reply or wait for you to finish it ;P

OpenStudy (anonymous):

Shadow you're in kindergarten right?

OpenStudy (shadowlegendx):

Yes

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

Or you can also draw like a 100 boxes and shade in 27...

OpenStudy (luigi0210):

Ain't nobody got fo dat batman

OpenStudy (shadowlegendx):

But uhh I don't know this percent equation that you speak of+

OpenStudy (luigi0210):

 Shade in a a fourth and a bit more >.>

OpenStudy (shadowlegendx):

Lui imma kill you

OpenStudy (shadowlegendx):

I just have to find my BaZoOkA

OpenStudy (luigi0210):

Fine, draw your 100 boxes >_>

OpenStudy (anonymous):

\[\frac{ Part }{ Whole } = \frac{ Percent }{ 100 }\] It's not really important for you at this moment :p

OpenStudy (anonymous):

Can kittens even do math?

OpenStudy (shadowlegendx):

1 + 1 = 11 so yeah

mathslover (mathslover):

Some great mathematicians here! ` 1 + 1 = 11`

OpenStudy (shadowlegendx):

Yesh @mathslover , behold I am the Great Math Kitty

OpenStudy (shadowlegendx):

So I guess I could probably just do a pie chart as it just asks for a model

OpenStudy (shadowlegendx):

Thanks for the suggestions guys!

OpenStudy (mathstudent55):

Here's a 10 by 10 grid. It contains a total of 100 unit-squares.